OpenStudy (anonymous):

Write the expression as a single natural logarithm. 3 ln 3 + 3 ln c

OpenStudy (anonymous):

factor out 3 3(ln 3 + ln c)

OpenStudy (anonymous):

can you do the rest?

OpenStudy (anonymous):

Use the following properrties: \[\large a \ln (b) = \ln (b^a)\] \[\ln a + \ln b = \ln (ab)\]

OpenStudy (anonymous):

You can write \[3 \ln(3) = \ln(3^3)\] \[3 \ln(c) = \ln(c^3)\] So you get \[\ln(3^3)+\ln(c^3)\] Now use the second property which I gave in my previous comment

OpenStudy (anonymous):

soo...\[\ln9c^3\]

OpenStudy (anonymous):

^that's right

OpenStudy (anonymous):

Wrong. 3^3 = 27

OpenStudy (anonymous):

you might wanna do a parenthesis though \[\large \ln (27c^3)\] and should be 27 lol

OpenStudy (anonymous):

:o ohh i see, hah oops thanksssss

OpenStudy (anonymous):

np :)